In a bowl throw 2.7 grams of aluminum in powder form and ignite it in air. the product obtained by combustion is dissolved with dilute hydrochloric acid. calculate how many grams of salt will be obtained from this case.
1
Expert's answer
2020-09-21T06:51:59-0400

4Al+3O2=2Al23O3

Al2O3+6HCl=2AlCl3+3H2O

according to the above reactions:

4Al=> 2Al2O3

2Al2O3=> 4AlCl3 that is,

1Al=> 1AlCl3

1. According to the above reaction, 27 g (1 mol) of Al produces 133.5 g (1 mol) of AlCl3. Find how much AlCl3 is formed from 2.7 g (0.1 mol) of Al:

27 gr Al — 133.5 gr AlCl3

2.7 gr Al — x gr AlCl3

x= 2.7*133.5/27= 13.35 gr

ANSWER: 13.35 gr AlCl3